OpenStudy (anonymous):

One number is 4 less than 3 times another number. If their sum is 16, find the numbers. What equations could be used to solve the problem if x represents a number?

OpenStudy (anonymous):

Lets say 'x' represents the "another number", then "one number is 4 less than 3 times another number is: \[3x-4\] So our 2 numbers are 'x' and '3x-4'. Their sum is 16, so now we have: \[x+ 3x-4 = 16 \Rightarrow 4x-4 = 16 \Rightarrow 4x = 20 \Rightarrow x = 5\] So the first number is 5, and the second number is 3(5)-4 = 11